OpenStudy (jamesj):

Here's a very nice question. Using the standard power series expansion of e^1, evaluate the following limit ...

OpenStudy (jamesj):

\[ \lim_{n \rightarrow \infty} n.\sin(2\pi.e.n!) \]

OpenStudy (amoodarya):